OpenStudy (anonymous):

Solve the system of equations. 3x + 3y = –6 –x + y = 0 A. x = –1; y = –1 B. x = 22; y = 7 C. x = 10; y = –3 D. x = 14; y = 5

OpenStudy (mertsj):

Plug each answer into both equations and choose the answer that makes both sentences true.

OpenStudy (anonymous):

okay. sometimes when i look at problems like these i think i have to go into a huge equation

OpenStudy (mertsj):

You could solve the system but it is simpler in this case to just plug and chug

OpenStudy (anonymous):

lol thanks

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

so it's A. right?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

thanks for helping me

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

here's a medal!

OpenStudy (mertsj):

ty

OpenStudy (anonymous):

anytime

OpenStudy (compassionate):

Hi, 3x + 3y = –6 –x + y = 0 There are two ways to solve this. By substitution or by elimination. The Elimination Method seems simpler in this case. We're going to multiply the bottom terms by -3, that way we can get the y terms to cancel. 3x + 3y = –6 -3(–x + y) = 0 3x + 3y = -6 3x - 3y = 6 Now our y-terms cancel and we have 6x = 0 (upon adding) Which turns into x = 0/6, which reduces to x = 0. Now plug in 0 for x in any of the equtions above. 3(0) - 3y = 6 -3y = 6 y = -2. So we have (0, -2) as our two terms. Plug it in to check your answer. 3(0) - 3(-2) = 6 0 + 6 = 6 6 = 6 Yes! The equation works! six is equal to six! This is known as the Elimination Method. I hope this was helpful and insightful! :) OpenStudy Ambassador: Compassionate

OpenStudy (anonymous):

yw! :D !

OpenStudy (anonymous):

thanks! if i could give medals to both of you i would!

OpenStudy (mertsj):

@Compassionate I'm afraid you made an error.

OpenStudy (mertsj):

3(0)=0 and -6+0= -6

OpenStudy (compassionate):

3x + 3y = –6 3(–x + y = 0) 3x + 3y = -6 -3x + 3y = 0 6y = -6 (Solving for x) y = -1 -x + y = 0 -x + -1 = 0 -x = 1 x = -1 I see my error @mertsj , I fixed it! Ignore the aforementioned post, dear. I messed up. Even we Ambies mess up sometimes! :) This is the correct solution, with work shown! Cheers ~

OpenStudy (compassionate):

It's the thought that counts! :)

OpenStudy (the_fizicx99):

Yesh it does ♥

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you guys all deserve one.

OpenStudy (anonymous):

just for trying to help me!

OpenStudy (the_fizicx99):

♦_♦

OpenStudy (anonymous):

:D !!!